OpenStudy (anonymous):

8. Essay question: show all work. Suzie’s flower shop estimates that supplies costs for one month will be given by the polynomial 2x^2 + 14x – 9. She estimates that utilities costs for one month is given by 12x – 3. Find and simplify an expression that would estimate Suzie’s total costs for one month.

jimthompson5910 (jim_thompson5910):

Total Cost = (Cost of Supplies) + (Cost of Utilities) Total Cost = (2x^2 + 14x – 9) + (12x – 3) Let me know what you get

OpenStudy (anonymous):

2x^2=26x-6

OpenStudy (anonymous):

not= + sign

jimthompson5910 (jim_thompson5910):

should be 2x^2+26x-12 note: -9+(-3) = -(9+3) = -(12) = -12

OpenStudy (anonymous):

ok i just forgot to put the - in there